OpenStudy (anonymous):

Find the equation of the tangent line to the graph of the given function at the point with the indicated x-coordinate f(x) = x + 3/x ; x=4

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

find f(4) and f'(4).... what u got for these values?

OpenStudy (anonymous):

f(4) = 19/4? not sure how to get f'(4)

OpenStudy (anonymous):

f(4) is correct... what is your f'(x) = ???

OpenStudy (anonymous):

write out the function this way so u can use the power rule: \(\large f(x) =x+3x^{-1} \)

OpenStudy (anonymous):

ok so the derivative is 1+(-3x)?

OpenStudy (anonymous):

not quite.... you forgot to subtract 1 from the exponent...

OpenStudy (anonymous):

-1+(-3x)

OpenStudy (anonymous):

i mean -3x^-2

OpenStudy (anonymous):

ok... that's better .. so f'(x) = ???

OpenStudy (anonymous):

1+(-3x^-2)

OpenStudy (anonymous):

yes... \(\large f'(x)=1-3x^{-2}=1-\frac{3}{x^2} \) so f'(4) = ????

OpenStudy (anonymous):

would i plug in 4 into that equation?

OpenStudy (anonymous):

yes.... f'(4) will give you the slope of the tangent line at x=4.

OpenStudy (anonymous):

1-3/(4)^2 = 13/16

OpenStudy (anonymous):

ok... good... so you have a point: (4, 19/4) and a slope: m = 13/16 looks like you can write the equation of the line using point-slope form...

OpenStudy (anonymous):

y = 13/16x+19/4?

OpenStudy (anonymous):

hang on... my answer is actually in point-slope form...

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

i got something different.... :( try again...

OpenStudy (anonymous):

y = 13/16x+4?

OpenStudy (anonymous):

this is what i got: point-slope form: \(\large y-\frac{19}{4}=\frac{13}{16}(x-4) \) slope-intercept form: \(\large y=\frac{13}{16}x+\frac{3}{2} \)

OpenStudy (anonymous):

ok thanks. the slope-intercept form was the one i was looking for

OpenStudy (anonymous):

thanks a lot for the help

OpenStudy (anonymous):

yw...:)